### v2

@@ 3.2 The Eight-Landmark Detection
Table  2 presents the results for the eight-landmark detection. Among the conditions, HRNet with shape subspace landmark refinement and an input image size of $384\times 288$ pixels achieved the best performance across all 8 landmarks. Also, the average MAE (Ave8, Ave2) using shape subspace landmark refinement was always equal to or less than that without the refinement, indicating that shape subspace effectively refines landmark locations.
Fig. 5 illustrates the detected landmarks in the eight-landmark detection using HRNet with shape subspace refinement, showing that the landmarks are accurately located.
Figure 8: The relative error when shape subspace (SS) landmark refinement is applied. The relative error by SS refinement represents the change in the mean absolute error (MAE) of all the landmarks when the shape subspace is applied in the 8-landmark detection task. Multiple peak detection rate indicates the ratio of test samples with multiple peaks in at least one heatmap. Each plot corresponds to a different input image size. As the multiple peak detection rate in the heatmap increases, the MAE decreases.
3.3 Point-based Segmentation
@@
Fig. 6 shows the point-based segmentation results using SAM2 with point prompts derived from HRNet-detected landmarks and shape subspace refinement. The green stars indicate positive prompts, the red stars denote negative prompts, and the blue areas represent segmented regions. This figure shows that the humerus and ulna are segmented with high fidelity.
4 Discussion
@@
First, heatmap-based methods outperformed other approaches in both results. Among them, HRNet achieved the highest performance. The ability of HRNet to capture local features while maintaining high-resolution representations can make it suitable for precise landmark detection in ultrasound images. In contrast, Transformer-based models like ViTPose typically require larger datasets for effective training. Given the limited dataset size, HRNet’s efficiency with smaller datasets appears to contribute to its superior performance.
@@
Third, in the eight-landmark model, shape subspace refined landmark positions. Fig. 7 displays a refinement example, with the LM3 heatmap having multiple peaks. Although the right peak has a higher value than the left peak in the LM3, the shape subspace refinement correctly selects the left peak that aligns with the ground truth position of landmark 3, as shown by the red point on Fig. 7(b).
Fig. 8 shows the relative error after applying shape subspace landmark refinement. As the ratio of multiple peaks in the heatmaps increases, the MAE decreases, demonstrating the effectiveness of shape subspace refinement. The refinement was more pronounced in ViTPose, likely because its broader context capture produces multiple peaks in heatmaps more frequently.
Finally, our method satisfy the accuracy needed for UCL injury diagnosis. While pitchers with UCL injuries show an average joint space increase of $1.2$ mm [2], our method achieved a $0.116$ mm average error using HRNet in two-landmark detection.
[…]
@@
Limitation First, the dataset was acquired using specific ultrasound equipment, which may not fully capture clinical variability. Second, despite the dataset comprising data from participants from diverse backgrounds, the model may have limited generalization to populations with different anatomies. Third, while shape subspace refines landmarks when multiple peaks exist, it may struggle when peaks are closely connected.
5 Conclusion
@@
In this paper, we introduced a novel ultrasound dataset consisting of 4,201 medial elbow images from 22 individuals, annotated with landmarks under the supervision of orthopedic experts. We evaluated joint space measurement methods on our dataset using multiple landmark detection models. The results showed that HRNet achieved a mean joint space measurement error of $0.116$ mm, demonstrating sufficient precision for detecting UCL injuries. In addition, we proposed a shape subspace landmark refinement method, which improved landmark localization performance. These suggest that our approach can facilitate the early diagnosis and intervention of UCL injuries by enabling automated joint space measurement. Furthermore, we demonstrated point-based bone segmentation of the humerus and ulna using detected landmarks as point prompts, highlighting a potential application of our dataset and methodology in musculoskeletal ultrasound analysis.
6 Declarations and Statements
6.1 Competing Interests
The authors declare no competing interests.
6.2 Funding
This study is supported by [Blinded for review].
6.3 Ethics Approval
This study complies with the 1964 Declaration of Helsinki. Approval was obtained by a committee of [Blinded for review].
6.4 Consent to Participate
Written consent was obtained from all participants.
6.5 Informed Consent and Consent for Publication
Not applicable. The ultrasound images used for this study do not contain personally identifiable information.
6.6 Data Availability Statement
The dataset is available at https://github.com/Akahori000/Ultrasound-Medial-Elbow-Dataset.

### v3

@@ VI-B The Eight-Landmark Detection
Table  II presents the results for the eight-landmark detection. Among the conditions, HRNet with SSR showed the best performance in the average MAE (Ave8, Ave2). Also, the average MAE using SSR was always equal to or less than that without SSR, indicating that shape subspace effectively refines landmark locations. Fig. 5 illustrates the detected landmarks in the eight-landmark detection using HRNet with SSR, which demonstrates that the landmarks are localized accurately. The fold-wise results for HRNet with SSR are summarized in Table III.
VI-C Point-based Segmentation
@@
Fig. 6(c) presents the point-based segmentation results of SAM2, using the detected landmarks from HRNet with SSR as point prompts. Green stars indicate positive prompts, red stars denote negative prompts, and blue areas represent segmented regions. This figure demonstrates that the humerus and ulna are segmented with high fidelity.
Figure 6: The point-based segmentation results using SAM2 with point prompts derived from the detected landmark positions. (a) Original images. (b) Ground truth of the humerus and ulna, shown in blue-green. (c) Segmented images of the humerus and ulna using SAM2 with point prompts derived from detected landmarks. The blue areas indicate the segmented regions, green stars mark the positive point prompts, and red stars denote the negative point prompts.
Figure 7: An example of landmark refinement by SSR. (a) The overlays of the output heatmaps on the input image. Multiple peaks are observed in the LM3 heatmap. (b) The localized landmarks after applying SSR. The detected landmarks 1–8 are shown in pink, light blue, red, yellow, green, blue, orange, and purple, respectively. In the LM3 heatmap, although the right peak exhibits higher intensity, SSR correctly selects the left peak, shown as the red point, which is closer to the ground truth.
VII Discussion
@@
First, the heatmap-based methods outperformed other approaches in both results. Among them, HRNet achieved the highest performance. The ability of HRNet to capture local features while maintaining high-resolution representations makes it particularly well-suited for precise landmark detection in ultrasound images. In contrast, Transformer-based models like ViTPose typically require larger datasets for effective training. Given the limited dataset size, HRNet’s efficiency with smaller datasets may have contributed to its superior performance.
@@
Figure 8: A pair of inference results on the medial elbow of a baseball player using our approach: (a) without stress, (b) with 2.5 kgf of applied valgus stress.
Figure 9: A representative failure case in which the model predicted the proximal end of the ulnar coronoid process to the left of the ground truth, likely due to a tissue-like structure visually connected to the ulna’s left edge.
Third, in the eight-landmark model, SSR effectively reduced landmark position errors. The degree of error reduction achieved by SSR varied across models: 0.025 mm for HRNet, 0.136 mm for ViTPose, and 0.161 mm for U-Net. This variation is attributed to differences in multiple peak detection rates in the output heatmaps, which were 12.1% for HRNet, 30.1% for ViTPose, and 66.2% for U-Net. These results indicate that SSR becomes more effective in improving landmark localization accuracy as the multiple peak detection rate increases. Fig. 7 shows an SSR example, with the LM3 heatmap having multiple peaks. In the LM3 heatmap, although the right peak has a higher value than the left peak, SSR correctly selects the left peak that aligns with the ground truth position of landmark 3, as shown by the red point in Fig. 7(b). The SSR inference required approximately 0.0006 seconds per image, and HRNet took 0.0128 seconds, resulting in a total of about 0.0134 seconds per image, making it suitable for real-time applications. This measurement was performed on a machine with an AMD Ryzen 9 5950X CPU and a single NVIDIA RTX 3090 GPU. Such rapid inference aligns with the broader trend of employing AI-based diagnosis [42], [43].
Finally, we evaluated whether our method meets the precision requirement for the diagnosis of UCL injuries. Pitchers with UCL injuries typically exhibit an average joint space increase of 1.2 mm under stressed conditions [2]. In comparison, our method achieved an average error of 0.125 mm using HRNet in two-landmark detection, which is substantially smaller than this pathological threshold. Also, to assess whether the 1.2 mm difference is distinguishable from measurement noise, we applied the 3-sigma rule. Let $\Delta$ denote the difference in joint space measurements between the stressed and unstressed conditions. Assuming equal variance under both conditions, the standard deviation of the difference is given by: $\sigma_{\Delta}=\sqrt{2}\,\sigma.$ This yields a limit of detection (LoD) defined as: $\mathrm{LoD}=3\sqrt{2}\,\sigma.$ Therefore, the measurement is considered distinguishable if: $\sigma\leq\frac{1.2}{3\sqrt{2}}\approx 0.283~\text{mm}.$ In five-fold cross-validation, four out of five folds in each split exhibited standard deviations ranging from 0.092 to 0.183 that are below the threshold 0.283 mm, while only one fold per split showed a higher deviation (up to 0.442) as shown in Table III (see Length values under Two-landmark detection). These results indicate that our method generally yields stable measurements, with increased variability observed only in one fold per split. This variability may be attributed to artifacts or imaging ambiguity, such as vague edges of the ulna and the presence of tissue-like structures that appear visually connected to the edge of the ulna, as illustrated in Fig. 9.
VII-A Limitations and Future Work
@@
First, the dataset was acquired using specific ultrasound equipment, which may not fully capture clinical variability, such as differences in machines and probes. Second, despite the dataset comprising data from participants from diverse backgrounds, the model may have limited generalization to populations with different anatomies. To address these limitations, future work could include collecting data across various devices and anatomical variations such as humerus avulsion cases and pediatric elbows, followed by model retraining to enhance robustness and clinical applicability. Third, while shape subspace refines landmarks when multiple peaks exist, it may struggle when peaks are closely connected over a few pixels or when no clear peak is present in the heatmap.
VIII Conclusion
@@
In this paper, we introduced a novel ultrasound dataset consisting of 4,109 medial elbow images from 22 individuals, annotated with landmarks under the supervision of orthopedic experts. Joint space measurement methods trained on our dataset were evaluated, with HRNet achieving a mean joint space measurement error of 0.125 mm, which demonstrated sufficient precision for detecting UCL injuries. In addition, we proposed a shape subspace refinement method, which effectively improved landmark localization performance. These results suggest that our approach can facilitate the early diagnosis and intervention of UCL injuries by enabling automated joint space measurement. Furthermore, the point-based bone segmentation of the humerus and ulna is demonstrated using detected landmarks as point prompts, highlighting a potential application of our dataset and methodology in musculoskeletal ultrasound analysis.
Acknowledgment
The authors appreciate Prof. Kazuhiro Fukui, Mr. Matheus Silva de Lima, and Mr. Santos Enoque Safrao for their guidance on the shape subspace implementation. Special thanks are extended to Prof. Xie Chun for his continuous support and insightful advice throughout this work. The authors would like to thank Prof. Yasukazu Totoki for his support as the principal investigator of the JSPS KAKENHI project (Grant Number 24K14397). Finally, gratitude is expressed to all the participants for their cooperation and consent.